OpenStudy (anonymous):

Arturo invests $2700 in a savings account that pays 9% interest, compounded quarterly. If there are no more other transactions,when will his balance reach $4550?

OpenStudy (anonymous):

Use the formula to set up the problem, then solve for the time.\[4550=2700(1+\frac{.09}{4})^{4t} \implies \frac{4550}{2700}=1.0225^{4t}\]There are a few steps after this, but you should be able to figure it out.

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (anonymous):

No sweat. Do math every day.